how large is an angle equal to eight times 15 degrees 45'?
Answer by stanbon(48522)

how large is an angle equal to eight times 15 degrees 45'?
000
45' is (3/4)minute
--------------------
15 degrees 45' = 15 3/4 degrees =
------------------------------------
8(15 3/4 degrees) = 120 + 6 = 126 degrees
===============================================
Cheers,
Stan H.